OVHCloud Bare Metal Cloud Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#6511 — Nas HA 10.16.100.91
Incident Report for Bare Metal Cloud
Resolved
After a problem of electrical maintenance, the whole leclerc on 10.16.100.91 was unavailable for a few minutes.
The problem was corrected.


The following paths were impacted :

zpool-000982
zpool-000986
zpool-000988
zpool-000995
Posted Mar 20, 2012 - 16:49 UTC